In a hypothetical reaction, 17.45 g chloric acid, HClO3, is reacted with 88.37 g tetraphosphorus decaoxide to form dichlorine heptoxide and an acid H2P4O9. Calculate the amount of the excess reactant left over.



1
Expert's answer
2021-09-10T00:42:33-0400

Ca(OH)2 + 2HCl = CaCl2 + 2H2O

m(CaCl2) real = 63.48 g

m(CaCl2) = 44.50 / 36.45 / 2 × 110.98 = 67.74g

Yield = 63.48/67.74 × 100% = 93.71%